OpenStudy (anonymous):

Consider a rational function p(x) with vertical asymptotes at x=3 and x=-3. The graph of p(x) shows a horizontal asymptote at y=4, and a y intercept at the point (0,-1). Write the function p(x)

OpenStudy (anonymous):

denominator must have zeros at \(3\) and \(-3\) so it could be \[(x+3)(x-3)=x^2-9\]

OpenStudy (anonymous):

horizontal asymptote at \(y=4\) means the numerator must have the same degree as the denominator and leading coefficient of 4, so it could be for example \(4x^2\)

OpenStudy (anonymous):

but since \(f(0)=-1\) it cannot be \(4x^2\) but it could be \(4x^2-1\)

OpenStudy (anonymous):

try \[p(x)=\frac{4x^2-1}{x^2-9}\] weird that they call it \(p\) since it is not a polynomial...